# Learning to Solve Multi-Objective Routing Problems on Multigraphs

### v1

@@ Appendix B Model Training
B.1 Curriculum Learning Details
@@
Additionally, we start by training GMS-DH for epochs (included in the first 100 epochs) on the XASY distribution with 20 nodes. This is because the selection-decoder relies on a reward signal from the routing-decoder. As such, we find that performance and rate of convergence improves if the routing-decoder is pre-trained to obtain a stable and accurate signal. We also find that stability increases if we freeze the parameters of the shared encoder with respect to the loss of the selection-decoder. That is, only the loss from the routing-decoder is utilized to calculate encoder gradients.
B.2 Training Times
@@
Appendix C Routing Problems and Distributions
@@ C.2 MOTSP
Problem Description: In the multi-objective traveling salesman problem the goal is to visit all nodes/customers and return to the starting position. The Euclidean version of this problem is widely treated in the learning-for-routing community, see e.g., Lin, Yang, and Zhang (2022) and Fan et al. (2025) in the multi-objective case. In our case, we sample two distances, corresponding to the two objectives, between each pair of customers. The goal is to design a tour that minimizes the sum of the distances. A tour can start from any node.
@@
Data Generation: Edge distances are sampled according to the EUC, TMAT and XASY distributions above.
@@
Model Details: There are no node features for this problem, hence both the GREAT-based and MatNet-based encoders have the two distances as edge features, while the latter uses zero-vectors as ”row” input.
@@
Evaluation Details: For all distributions, we utilize , and as reference for HV calculation for , and nodes respectively. In the zero-shot experiments, we have as reference for 150 nodes and for 200 nodes.
C.3 MOCVRP
@@
Problem Description: In the multi-objective capacitated vehicle routing problem, a truck starts at a depot and must pick up goods from a specified number of customers. The vehicle has a fixed capacity while each customer has a demand. When the capacity is full, the vehicle must return to the depot. As many vehicles as required can be utilized to meet the demand of all customers.
@@
Similarly to the TSP, this problem has been widely treated in the Euclidean setting using learning-based methods. As is common in the multi-objective setting (Lin, Yang, and Zhang 2022), we assume one objective is the traveled distance, while the other is the longest tour for a single vehicle (makespan). As such, each node-pair is only associated with one distance.
@@
Data Generation: We assume that the capacity of the truck is 30, 40 and 50 for 20, 50 and 100 nodes respectively, while the demands are sampled uniformly from the set . This is the same setting as in e.g., Kool, van Hoof, and Welling (2018) and Lin, Yang, and Zhang (2022). Edge distances are sampled according to the EUC, TMAT and XASY distributions.
@@
Model Details: In the decoding, we augment the routing context (represented by the query ) with the current load of the vehicle. For the GREAT-based encoders, we let the features of an edge be its distance as well as the demand of the customer at the end. On the other hand, the MatNet-based encoders only utilize the distance as an edge-feature while the demand is used as a node feature.
@@
Evaluation Details: For all distributions, we utilize , and as reference for HV calculation for the three considered instance sizes.
C.4 MGMOTSP
@@
Problem Description: As for the MOTSP, the goal of the multigraph MOTSP is to visit all nodes and return to the start. The only difference is that the underlying structure is a multigraph. Similarly, we define two distances for each node pair and these are associated with the objectives of the problem.
@@
While this variant of the TSP on a multigraph has not been treated before to our knowledge (neither using learning nor other methods), some stochastic single-objective variants have been proposed in e.g., Tadei, Perboli, and Perfetti (2017) and Maggioni, Perboli, and Tadei (2014). Moreover, we believe that the MGMOTSP is a suitable benchmark problem for our solvers, as it is simple and easy to find good baselines for.
Data Generation: We sample distances between the nodes according to the FLEX and FIX distributions.
@@
Model Details: As the MOTSP, the MGMOTSP lacks edge features. Hence, we utilize the same setup for the models as in the MOTSP. Moreover, for the MatNet-based model, we sparsify the underlying multigraph given each preference before it is fed into the encoder. This is done by removing all edges with sub-optimal LS cost compared to a parallel edge. To make this pruning strategy exact, we train MBM using linearly scalarized reward instead of the Chebyshev reward.
@@
Evaluation Details: For the FLEX distribution, we utilize , and as reference for instance sizes 20, 50 and 100, regardless of . In the zero-shot experiments we set the references to and for FLEX with 150 and 200 nodes. For FIX, the reference is given by , where is the node count.
C.5 MGMOTSPTW
@@
Problem Description: The goal in MGMOTSPTW is to create a tour that visits all customers and returns to the starting position. However, in this case we assume there is a fixed depot and that each customer/node is associated with a time-window. Preferably, the vehicle should arrive within the time-window for each customer. As such, we let one objective be the number of violated time-windows while the other is the traveled distance. Accordingly, each edge connecting two locations is associated with a travel time and a distance.
@@
Multi-objective problems with similar objectives are fairly widespread in literature (Legrand et al. 2025), although not on multigraphs. In the multigraph framework, this problem is reminiscent of the vehicle routing problem with time-windows in Ben Ticha et al. (2017), but with a second objective that is the soft variant of the time-window constraint.
@@
Data Generation: Distances and travel times are sampled according to the FLEX and FIX distributions. Additionally, we sample time-windows according to the ”medium” distribution of Chen et al. (2024) and Bi et al. (2024). In this setting, time-windows are sampled uniformly and independently, which results in difficult instances where there is a true trade-off between optimal paths and paths satisfying the time-window constraint. We also argue that this premise is simple for initial evaluation of our models, especially in comparison to e.g., the ”hard” instances of Chen et al. (2024), which tests out-of-distribution generalization.
@@
Model Details: In the decoder, we enhance the routing context with an embedding of the current time. In the encoder, the node features in this case are the starting time and end time of the service, while the edge features are distance and travel time. Consequently, we augment the input features of GREAT so that an edge is associated with its distance, travel time and the time-window of the ending node. As usual for MatNet, the node features are used as ”row” inputs while the edge features are used in the distance matrix.
Note that in this case, pruning of the multigraph for MatNet is not obvious. For instance, removing an edge with a high travel time but low distance might not be optimal if traversing that edge would result in no violated time-windows anyway. Consequently, we apply the same simple pruning method as for the MGMOTSP. That is, we form the cost of each edge using a linear combination of its travel time and distance, then we remove all edges apart from the best one. We also choose the linear reward instead of the Chebyshev reward for MBM.
Evaluation Details: In this case, we let the reference value corresponding to the time-window violation be , while the reference value corresponding to the distance is the same as in the MGMOTSP case.
Appendix D Setup for MOEAS
As a base for the implementation of NSGA-II, NSGA-III and MOEA/D, we utilized Pymoo222https://pymoo.org/ and as a base for MOGLS, we utilized GeneticLocalSearch_TSP333https://github.com/kevin031060/Genetic˙Local˙Search˙TSP. For the former three, we limit the number of iterations to while the latter is limited to . Additionally, local search strategies is a necessity for efficient usage of genetic algorithms in combinatorial optimization. For the MOTSP, we utilized the 2-opt heuristic of Jaszkiewicz (2002), while we utilized the problem-specific local search and chromosome decoding strategy of Lacomme, Prins, and Sevaux (2006) for the MOCVRP.
However, in both cases we find that by accepting the best local move instead of just an improving one, the performance on large asymmetric problems improves substantially. As this affects the runtime negatively (due to the need to search through all moves), we limit the number of local moves in each search depending on the problem size. Additionally, for the algorithms not requiring preference weights for each individual, we must calculate such weights to compare the improvements. In the bi-objective case, this is done using the strategy of Lacomme, Prins, and Sevaux (2006) according to
where is the solution being evaluated and , are minimum and maximum values for the respective objectives.
@@ Appendix D Setup for MOEAS
For the mutation operator, we randomly either change one edge or reverse a segment of the tour. For the crossover, we utilize an edge recombination operator. Regarding the local search, we use a variant of 2-opt where we again accept the best local move. A local move is defined by a standard 2-opt move with the nodes, together with new edge selections given the changed node order. This edge selection is found by choosing the edge with the lowest LS cost between nodes that were previously not neighbors or have changed order.
We want to stress that the focus in this study is not to design an evolutionary algorithms, we simply want a baseline to compare against. We find that the proposed approach is somewhat efficient for small node counts, but performs quite badly for large graphs.
Appendix E Complementary Numerical Results
@@
E.1 Problems on Simple Graphs
Table 6 and 7 show results for the MOTSP and MOCVRP respectively. Apart from the benchmarks in the main text, we present the performance of three simple heuristics: Nearest Neighbor (NN), Nearest Insertion (NI) and Farthest Insertion (FI). As LKH and OR-tools, these utilize the linear scalarization to transform the MO problem into single-objective subproblems. Additionally, we show results on Euclidean instances (EUC). For these, there are many existing models in literature and thus we replace MBM with the Conditional Neural Heuristic (CNH) of Fan et al. (2025). This model is augmented with the Euclidean augmentation of Lin, Yang, and Zhang (2022).
On the EUC distribution, our models are generally slightly worse than CNH. This is to be expected though, as a common pattern in previous studies is that architectures built for directed graph inputs perform worse on EUC than those specialized for coordinate inputs (Lischka et al. 2024; Kwon et al. 2021).
@@ E.1 Problems on Simple Graphs
E.2 Multigraph Problems
We display additional results for the MGMOTSP and MGMOTSPTW in Tables 8-10. As for the MOTSP, we include NN, NI and FI for MGMOTSP. In the case of MGMOTSPTW, we also include results for LKH and OR-tools when treating the problem as the MGMOTSP. That is, we neglect the time-windows and treat the time duration of each edge exactly as the distance. We call these methods implicit LKH and implicit OR-tools, as they implicitly minimize the time-window violations by reducing the time the tour takes.
@@ E.2 Multigraph Problems
| MOGLS | 0.71 | 16.29% | (1.9h) | 0.49 | 45.84% | (29h) | 0.43 | 54.91% | (18h) |
| NSGA-II | 0.81 | 5.06% | (1.3h) | 0.74 | 17.62% | (19h) | 0.64 | 32.11% | (15h) |
| NSGA-III | 0.80 | 5.82% | (1.3h) | 0.74 | 18.48% | (19h) | 0.63 | 33.16% | (14h) |
| MOEA/D | 0.80 | 6.02% | (1.4h) | 0.74 | 17.50% | (18h) | 0.66 | 30.28% | (12h) |
@@
| MOGLS | 0.74 | 13.68% | (1.9h) | 0.65 | 29.64% | (29h) | 0.52 | 45.00% | (19h) |
| NSGA-II | 0.81 | 5.64% | (1.3h) | 0.77 | 16.56% | (19h) | 0.69 | 27.94% | (15h) |
| NSGA-III | 0.80 | 6.85% | (1.3h) | 0.75 | 18.79% | (19h) | 0.66 | 30.49% | (15h) |
| MOEA/D | 0.80 | 6.20% | (1.4h) | 0.78 | 15.42% | (18h) | 0.69 | 27.63% | (12h) |
@@
E.3 Generalization to Larger Instances
In Tables 11-12, we report the generalization performance zero-shot for our methods in the MOTSP and MGMOTSP cases. It can be seen that both models retain competitive performance. In particular, GMS-EB often performs close to LKH while being faster.

### v2

@@ Appendix C Model Training
C.1 Curriculum Learning Details
@@
Additionally, we start by training GMS-DH for epochs (included in the first 100 epochs) on the XASY distribution with 20 nodes. This is because the selection-decoder relies on a reward signal from the routing-decoder. As such, we find that performance and rate of convergence improves if the routing-decoder is pre-trained on simple graphs to obtain a stable and accurate signal. We also find that stability increases if we freeze the parameters of the shared encoder with respect to the loss of the selection-decoder. That is, only the loss from the routing-decoder is utilized to calculate encoder gradients.
C.2 Training Times
@@
Appendix D Routing Problems and Distributions
@@ D.2 MOTSP
In the multi-objective traveling salesman problem the goal is to visit all nodes/customers and return to the starting position. The Euclidean version of this problem has been widely treated in the learning-for-routing community, e.g., by Lin, Yang, and Zhang (2022) and Fan et al. (2025). In our case, we sample two distances, corresponding to the two objectives, between each pair of customers. The goal is to design a tour that minimizes the sum of the distances. A tour can start from any node.
Data Generation
@@
Edge distances are sampled according to the EUC, TMAT and XASY distributions above.
Model Details
@@
There are no node features for this problem, hence both the GREAT-based and MatNet-based encoders have the two distances as edge features, while the latter uses zero-vectors as ”row” input.
Evaluation Details
@@
For all distributions, we utilize , and as reference for HV calculation for , and nodes respectively. In the zero-shot experiments, we have as reference for 150 nodes and for 200 nodes.
D.3 MOCVRP
Problem Description
@@ D.3 MOCVRP
In the multi-objective capacitated vehicle routing problem, a truck starts at a depot and must pick up goods from a specified number of customers. The vehicle has a fixed capacity while each customer has a demand. When the capacity is full, the vehicle must return to the depot. As many vehicles as required can be utilized to meet the demand of all customers.
@@
Similarly to the TSP, this problem has been widely treated in the Euclidean setting using learning-based methods. As is common in the multi-objective setting (Lin, Yang, and Zhang 2022), we assume one objective is the traveled distance, while the other is the longest tour for a single vehicle (makespan). As such, each node pair is only associated with one distance.
Data Generation
@@
We assume that the capacity of the truck is 30, 40 and 50 for 20, 50 and 100 nodes respectively, while the demands are sampled uniformly from the set . This is the same setting as in e.g., Kool, van Hoof, and Welling (2018) and Lin, Yang, and Zhang (2022). Edge distances are sampled according to the EUC, TMAT and XASY distributions.
Model Details
@@
In the decoding, we augment the routing context (represented by the query ) with the current load of the vehicle. For the GREAT-based encoders, we let the features of an edge be its distance as well as the demand of the customer at the end. On the other hand, the MatNet-based encoders only utilize the distance as an edge-feature while the demand is used as a node feature.
Evaluation Details
@@
For all distributions, we utilize , and as reference for HV calculation for the three considered instance sizes.
D.4 MGMOTSP
Problem Description
@@ D.4 MGMOTSP
As for the MOTSP, the goal of the multigraph MOTSP is to visit all nodes and return to the start. The only difference is that the underlying structure is a multigraph. Similarly, we define two distances for each node pair and these are associated with the objectives of the problem.
@@
While this variant of the TSP on a multigraph has not been treated before to our knowledge (neither using learning nor other methods), some stochastic single-objective variants have been proposed by e.g., Tadei, Perboli, and Perfetti (2017) and Maggioni, Perboli, and Tadei (2014). Moreover, we believe that the MGMOTSP is a suitable benchmark problem for our solvers, as it is simple and easy to find good baselines for.
Data Generation
We sample distances between the nodes according to the FLEX and FIX distributions.
Model Details
@@
As the MOTSP, the MGMOTSP lacks node features. Hence, we utilize the same setup for the models as in the MOTSP. Moreover, for the MatNet-based model, we sparsify the underlying multigraph given each preference before it is fed into the encoder. This is done by removing all edges with suboptimal linearly scalarized cost compared to a parallel edge. To make this pruning strategy exact, we train MBM using linear reward instead of the Chebyshev reward.
Evaluation Details
@@
For the FLEX distribution, we utilize , and as reference for instance sizes 20, 50 and 100, regardless of . In the zero-shot experiments we set the references to and for FLEX with 150 and 200 nodes. For FIX, the reference is given by , where is the node count.
D.5 MGMOTSPTW
Problem Description
@@ D.5 MGMOTSPTW
The goal in MGMOTSPTW is to create a tour that visits all customers and returns to the starting position. However, in this case we assume there is a fixed depot and that each customer/node is associated with a time-window. Preferably, the vehicle should arrive within the time-window for each customer. As such, we let one objective be the number of violated time-windows while the other is the traveled distance. Accordingly, each edge connecting two locations is associated with a travel time and a distance.
@@
Multi-objective problems with similar objectives are fairly widespread in literature (Legrand et al. 2025), although not on multigraphs. In the multigraph framework, this problem is reminiscent of the vehicle routing problem with time-windows of Ben Ticha et al. (2017), but with a second objective that is the soft variant of the time-window constraint.
Data Generation
@@
Distances and travel times are sampled according to the FLEX and FIX distributions. Additionally, we sample time-windows according to the ”medium” distribution of Chen et al. (2024) and Bi et al. (2024). In this setting, time-windows are sampled uniformly and independently, which results in difficult instances where there is a true trade-off between optimal paths and paths satisfying the time-window constraint. We also argue that this premise is simple for initial evaluation of our models, especially in comparison to e.g., the ”hard” instances of Chen et al. (2024), which test out-of-distribution generalization.
@@
Model Details
In the decoder, we enhance the routing context with an embedding of the current time. In the encoder, the node features in this case are the starting time and end time of the service, while the edge features are distance and travel time. Consequently, we augment the input features of GREAT so that an edge is associated with its distance, travel time and the time-window of the ending node. As usual for MatNet, the node features are used as ”row” inputs while the edge features are used in the distance matrix.
Note that in this case, pruning of the multigraph for MatNet is not obvious. For instance, removing an edge with a high travel time but low distance might not be optimal if traversing that edge would result in no violated time-windows anyway. For simplicity, we apply the same simple pruning method as for the MGMOTSP. That is, we form the scalarized cost of each edge using a linear combination of its travel time and distance, then we remove all edges apart from the best one. We also choose the linear reward instead of the Chebyshev reward for MBM.
Evaluation Details
In this case, we let the reference value corresponding to the time-window violation be , while the reference value corresponding to the distance is the same as in the MGMOTSP case.
Appendix E Setup for MOEAs
As a base for the implementation of NSGA-II, NSGA-III and MOEA/D, we utilized Pymoo222https://pymoo.org/ (Blank and Deb 2020) and as a base for MOGLS, we utilized GeneticLocalSearch_TSP333https://github.com/kevin031060/Genetic˙Local˙Search˙TSP (Li, Zhang, and Wang 2021). For the former three, we limit the number of iterations to while the latter is limited to . Additionally, local search strategies are a necessity for efficient usage of genetic algorithms in combinatorial optimization. For the MOTSP, we utilized the 2-opt heuristic of Jaszkiewicz (2002), while we utilized the problem-specific local search and chromosome decoding strategy of Lacomme, Prins, and Sevaux (2006) for the MOCVRP.
However, in both cases we find that by accepting the best local move instead of just an improving one, the performance on large asymmetric problems improves substantially. As this affects the runtime negatively (due to the need to search through all moves), we limit the number of local moves in each search depending on the problem size. Additionally, for the algorithms not requiring preference weights for each individual (NSGA-II and NSGA-III), we must calculate such weights to compare the improvements. In the bi-objective case, this is done using the strategy of Lacomme, Prins, and Sevaux (2006) according to
where is the solution being improved and , are minimum and maximum values for the respective objectives.
@@ Appendix E Setup for MOEAs
For the mutation operator, we randomly either change one edge or reverse a segment of the tour. For the crossover, we utilize an edge recombination operator. Regarding the local search, we use a variant of 2-opt where we again accept the best local move. A local move is defined by a standard 2-opt move with the nodes, together with new edge selections given the changed node order. This edge selection is found by choosing the edge with the lowest linearly scalarized cost between nodes that were previously not neighbors or have changed order.
We want to stress that the focus in this study is not to design evolutionary algorithms, we simply want baselines to compare against. We find that the proposed approach is somewhat efficient for small node counts, but performs quite badly for large graphs.
Appendix F Ablation Study
Here, we present additional results obtained by replacing the selection-decoder in GMS-DH with a simple pruning function that discards edges with suboptimal linearly scalarized cost. As shown in Section 5.4, this approach performs rather poorly on the MGMOTSPTW. However, it is more reasonable to assume that it will yield good results for the MGMOTSP.
To ensure a fair comparison, we evaluate both learned pruning and simple pruning using linearly scalarized rewards and Chebyshev rewards. In the linear case, simple pruning is optimal by Proposition 1. Moreover, the learned pruning model is trained using edge selections per instance. Accordingly, we use 4 more POMO samples when training the model with simple pruning to match the training budget.
Ablation results are shown in Table 5. For the MGMOTSP, simple and learned pruning perform quite similarly (although learned pruning shows a slight advantage under Chebyshev rewards for both distributions). Indeed, these results illustrate that it can be a smart approach to replace the selection-decoder in GMS-DH if a sound heuristic or exact pruning method is available. However, pruning after the encoder clearly outperforms pre-pruning (as used in GMS-DH PP), both in terms of solution quality and runtime, since pre-pruning requires reencoding the graph for each preference.
FLEX2-50 | FIX2-50 | |||
HV | Gap | HV | Gap | |
GMS-DH Learned (Ch. Reward) | 0.89 | 1.84% | 0.91 | 1.73% |
GMS-DH Learned (Lin. Reward) | 0.88 | 2.74% | 0.91 | 1.66% |
GMS-DH Simple (Ch. Reward) | 0.88 | 2.00% | 0.90 | 1.82% |
GMS-DH Simple (Lin. Reward) | 0.88 | 2.65% | 0.91 | 1.76% |
GMS-DH PP (Lin. Reward) | 0.77 | 14.24% | 0.82 | 10.51% |
Appendix G Complementary Numerical Results
@@
G.1 Problems on Simple Graphs
Table 6 and 7 show results for the MOTSP and MOCVRP respectively. Apart from the benchmarks in the main text, we present the performance of three simple greedy heuristics on the MOTSP: Nearest Neighbor (NN), Nearest Insertion (NI) and Farthest Insertion (FI). Like LKH and OR-tools, these utilize the linear scalarization to transform the MO problem into single-objective subproblems. Additionally, we show results on Euclidean instances (EUC). For these, there are many existing models in literature and thus we replace MBM with the Conditional Neural Heuristic (CNH) of Fan et al. (2025). This model is augmented with the Euclidean augmentation of Lin, Yang, and Zhang (2022).
On the EUC distribution, our models are generally slightly worse than CNH. This is to be expected though, as a common pattern in previous studies is that architectures built for directed graph inputs perform worse on EUC than those specialized for coordinate inputs (Lischka et al. 2025; Kwon et al. 2021).
@@ G.1 Problems on Simple Graphs
G.2 Multigraph Problems
We display additional results for the MGMOTSP and MGMOTSPTW in Tables 8-11. As for the MOTSP, we include NN, NI and FI for MGMOTSP. In the case of MGMOTSPTW, we also include results for LKH and OR-tools when treating the problem as the MGMOTSP. That is, we neglect the time-windows and treat the time duration of each edge exactly as the distance. We call these methods implicit LKH and implicit OR-tools, as they implicitly minimize the time-window violations by reducing the time the tour takes. Additionally, for the MGMOTSP, we show results for FLEX and FIX, i.e., when there are 5 parallel edges between each node pair.
@@ G.2 Multigraph Problems
MOGLS | 0.71 | 16.29% | (1.9h) | 0.49 | 45.84% | (29h) | 0.34 | 64.38% | (129h) |
NSGA-II | 0.81 | 5.06% | (1.3h) | 0.74 | 17.62% | (19h) | 0.64 | 31.98% | (103h) |
NSGA-III | 0.80 | 5.82% | (1.3h) | 0.74 | 18.48% | (19h) | 0.63 | 33.01% | (104h) |
MOEA/D | 0.80 | 6.02% | (1.4h) | 0.74 | 17.50% | (18h) | 0.68 | 28.46% | (99h) |
@@
FLEX5-20 | FLEX5-50 | FLEX5-100 | |||||||
LKH | 0.93 | 0.00% | (1.3m) | 0.95 | 0.00% | (6.4m) | 0.97 | 0.00% | (30m) |
OR Tools | 0.91 | 1.51% | (2.9m) | 0.93 | 2.01% | (25m) | 0.96 | 1.60% | (1.8h) |
NN | 0.86 | 6.78% | (1.8s) | 0.90 | 5.24% | (7.2s) | 0.94 | 3.33% | (28s) |
NI | 0.88 | 5.48% | (9.5s) | 0.89 | 6.41% | (1.7m) | 0.92 | 5.45% | (12m) |
FI | 0.87 | 5.89% | (11s) | 0.89 | 7.03% | (2.0m) | 0.92 | 5.89% | (14m) |
MOGLS | 0.84 | 9.71% | (2.4h) | 0.69 | 27.67% | (52h) | 0.57 | 41.21% | (222h) |
NSGA-II | 0.88 | 5.19% | (1.4h) | 0.80 | 16.61% | (31h) | 0.71 | 26.79% | (215h) |
NSGA-III | 0.87 | 5.85% | (1.4h) | 0.79 | 17.48% | (31h) | 0.70 | 27.98% | (215h) |
MOEA/D | 0.88 | 5.07% | (1.5h) | 0.82 | 13.49% | (30h) | 0.76 | 22.05% | (210h) |
MBM | 0.90 | 2.38% | (4.4s) | 0.93 | 2.48% | (13s) | 0.96 | 1.71% | (45s) |
MBM (aug) | 0.92 | 1.26% | (32s) | 0.93 | 2.12% | (1.7m) | 0.96 | 1.44% | (5.9m) |
GMS-EB | 0.92 | 0.87% | (8.1s) | 0.94 | 1.08% | (2.3m) | 0.96 | 1.11% | (23m) |
GMS-DH | 0.92 | 1.03% | (3.5s) | 0.94 | 1.64% | (16s) | 0.96 | 1.66% | (1.3m) |
GMS-EB (aug) | 0.92 | 0.59% | (1.1m) | 0.95 | 0.86% | (18m) | 0.96 | 0.96% | (3.0h) |
GMS-DH (aug) | 0.92 | 0.64% | (23s) | 0.94 | 1.31% | (2.2m) | 0.96 | 1.45% | (9.9m) |
@@
MOGLS | 0.74 | 13.68% | (1.9h) | 0.65 | 29.64% | (29h) | 0.54 | 42.73% | (122h) |
NSGA-II | 0.81 | 5.64% | (1.3h) | 0.77 | 16.56% | (19h) | 0.69 | 27.86% | (96h) |
NSGA-III | 0.80 | 6.85% | (1.3h) | 0.75 | 18.79% | (19h) | 0.66 | 30.17% | (95h) |
MOEA/D | 0.80 | 6.20% | (1.4h) | 0.78 | 15.42% | (18h) | 0.72 | 24.69% | (91h) |
@@
GMS-DH Simple | 0.76 | 8.18% | (3.3s) | 0.83 | 7.01% | (13s) |
FIX2-20 | FIX2-50 | |||||
HV | Gap | Time | HV | Gap | Time | |
Implicit LKH | 0.28 | 67.24% | (1.3m) | 0.15 | 84.24% | (6.5m) |
Implicit OR Tools | 0.25 | 71.02% | (3.4m) | 0.14 | 85.49% | (26m) |
MOGLS | 0.62 | 28.85% | (2.8h) | 0.41 | 55.72% | (21h) |
NSGA-II | 0.73 | 15.93% | (2.7h) | 0.56 | 39.66% | (35h) |
NSGA-III | 0.71 | 18.88% | (2.5h) | 0.53 | 43.42% | (35h) |
MOEA/D | 0.74 | 14.99% | (2.8h) | 0.60 | 35.73% | (36h) |
MBM | 0.53 | 38.64% | (4.9s) | 0.64 | 31.18% | (14s) |
MBM (aug) | 0.54 | 38.10% | (35s) | 0.66 | 28.58% | (1.9m) |
GMS-EB | 0.86 | 1.17% | (4.3s) | 0.93 | 0.55% | (59s) |
GMS-DH | 0.85 | 2.47% | (3.4s) | 0.91 | 1.74% | (12s) |
GMS-EB (aug) | 0.87 | 0.00% | (31s) | 0.93 | 0.00% | (7.9m) |
GMS-DH (aug) | 0.87 | 0.46% | (20s) | 0.92 | 0.78% | (1.7m) |
GMS-DH Simple | 0.76 | 12.07% | (3.2s) | 0.85 | 8.93% | (12s) |
G.3 Generalization to Larger Instances
In Tables 12-14, we report more generalization results for our methods in the MOTSP and MGMOTSP cases. Notably, even when introducing significantly more edges by testing on FLEX and FIX, our methods remain competitive zero-shot.